# Efficient Navigation in Unknown Indoor Environments with Vision-Language Models

### v1

@@ Abstract
We present a novel high-level planning framework that leverages vision-language models (VLMs) to improve autonomous navigation in unknown indoor environments with many dead ends. Traditional exploration methods often take inefficient routes due to limited global reasoning and reliance on local heuristics. In contrast, our approach enables a VLM to reason directly about an occupancy map in a zero-shot manner, selecting subgoals that are likely to lead to more efficient paths. At each planning step, we convert a 3D occupancy grid into a partial 2D map of the environment, and generate candidate subgoals. Each subgoal is then evaluated and ranked against other candidates by the model. We integrate this planning scheme into DYNUS [1], a state-of-the-art trajectory planner, and demonstrate improved navigation efficiency in simulation. The VLM infers structural patterns (e.g., rooms, corridors) from incomplete maps and balances the need to make progress toward a goal against the risk of entering unknown space. This reduces common greedy failures (e.g., detouring into small rooms) and achieves about 10% shorter paths on average.
Supplementary Material
@@ I Introduction
Autonomous mobile robots can be applied in numerous domains such as package delivery and search-and-rescue. In many of these applications, the robots need to operate in unknown environments where prior maps are not available. When navigating such spaces, traditional approaches such as DYNUS [1], often take inefficient routes by entering dead ends like small rooms, due to their lack of higher-level reasoning abilities. For instance, DYNUS simply projects the goal location onto a local map around the robot, and generates paths without global reasoning.
In recent years, many works have been proposed to improve the planning capabilities of robotic systems by employing large language models (LLMs) [2, 3, 4], including in frontier exploration schemes [5] and object-goal navigation [6]. We aim to leverage the reasoning competencies of multi-modal LLMs, in particular VLMs, to achieve efficient routing in unknown indoor environments that typically contain many impasses. The novelty of our approach lies in allowing VLMs to directly reason about an incomplete map, not RGB images. At each step, we convert the 3D occupancy grid into a voxelized 2D map, generate candidate subgoals, and ask the VLM to score directions in the context of the partial map (e.g., infer likely rooms/corridors from partial structure despite occlusions). We then aggregate multiple VLM queries, rank candidates by expected path efficiency, and select the top subgoal.
@@ III-A Problem Formulation
Our goal is to enable autonomous robots to traverse unknown indoor spaces without entering dead ends, using only occupancy data and no prior maps. At each decision step, the high-level planner receives (1) the terminal goal position and (2) a 3D occupancy grid that is built online. The planner outputs a navigation subgoal that is likely to result in the most direct possible path to the goal by reducing detours and backtracking.
@@ III-B Representing the Environment
The only information we expect to receive from the robot is a 3D occupancy grid of the environment. Note that we do not rely on RGB images, which require good lighting conditions, enabling our system to work in low-light conditions. However, the vision-language model cannot parse and understand raw sensor data or a 3D occupancy map. To obtain useful navigation guidance from the VLM, we ground the VLM with a map-like representation: a projected 2D occupancy map annotated with the robot pose, the goal, and candidate subgoals. This makes spatial relations explicit so the VLM can reason about where to go next. An example of such a representation is shown in Fig. 2.
@@ III-C Converting a 3D Occupancy Grid to a 2D Map
We keep a voxel map updated online. The map encodes unknown/free/occupied cells and overlays the robot pose, the goal (or its projection), and candidate subgoals. A cell is marked as free only if there is sufficient $z$-axis clearance for the robot and is reachable. We pass the map to the model either as a string or as an image. This can be achieved by associating each grid cell with a character or a color, respectively.
Because the input grid resolution is fixed, we specify the desired size of the map in both meters and grid cells, and down-sample the grid accordingly. To get the best possible approximation:
@@ III-F Selecting a Subgoal
The median has the advantage that it is robust to outliers, and does not blend different probabilities. In taking the mean of two valid but different solutions, for instance, one may not necessarily end up with a good answer. On top of that, we compute the median absolute deviation (MAD) for each belief as an uncertainty measure:
@@ V Conclusions
This paper presents a VLM-based high-level planning framework that selects navigation subgoals at frontiers between known and unknown space in indoor environments. By providing occupancy maps as input, the model can reason about basic geometric structure (e.g., rooms and corridors). The method assumes minimal prior knowledge about the environment or sensors and integrates cleanly with existing autonomy stacks.
In partially observed settings, avoiding dead ends requires making decisions under uncertainty. Our simulations show that the proposed approach avoids most dead ends and yields shorter paths on average than the baseline in DYNUS [1]. We also observe that VLMs can infer structural patterns from incomplete maps and make reasonable trade-offs when uncertainty is high.
The main limitation is model latency, which constrains real-time deployment. We expect this to improve with faster hardware and models. Future work includes adding semantic inputs to further inform decisions and support richer trade-offs.
Appendix
V-A Developer Prompt
@@ 1. Candidate Move Evaluation
- Enclosure Check:
@@
Developer Prompt (2/2) - Progress Toward Goal:
@@
Candidate Move Comparison
@@ V-B Example Responses
    ’This candidate is to the far north, up an extended corridor and within a semi-enclosed room. There is no visible path southward toward the projected goal. The route is circuitous and likely inefficient, with higher enclosure risk.’
    ’Southernmost exit from central area, leading south and slightly east. Appears to open onto a possible corridor or hallway extending south, roughly matching the direction of the goal. No signs of imminent enclosure.’
@@
    ’Northeast frontier, but accessing it requires backtracking through already-explored space. No clear indication it leads toward the goal. Likely a dead-end room given the local geometry and surrounding walls.’
    ’This candidate is located at the top-left part of the explored area, apparently in a cul-de-sac (enclosed by walls on north and west sides). No visible path toward the goal projection to the far southeast. High risk of being a dead end.’
@@
    ‘This candidate is north of the robot and to the east, leading into a corridor in a section with more open unexplored area. There’s a good chance it opens up, but it’s not the closest to the direction of the goal projection.’
    ’This is the southernmost candidate, ending the current visible main corridor. It is the frontier closest to the projected goal position, opens directly onto unknown space, and is highly promising for discovering the route to the goal. However, possible enclosure just south, so there is still risk.’
@@
    ’This is to the south of the current room and could be a passageway, but still not fully aligned with the more southeast position of the goal. Low enclosure risk, but the area seems more like an open room than a corridor.’
    ’This move goes south, away from the central room, and may be a corridor leading farther south. Though not directly leading southeast, it’s the only exit on the lower half; if this is a corridor, it could bend toward the east, following the general direction of the goal projection. Medium uncertainty.’
V-C grid-world Simulation Tool
See Fig. 4.
Figure 4: View of grid-world, a simulation tool we built to study the navigation capabilities of LLMs/VLMs. The 2D map on the left shows the robot (red) traversing a corridor in an effort to reach the final goal (green) in the top right. The squares in the middle and on the right display the normalized median belief and the median absolute deviation (MAD) of the belief, respectively, for each candidate subgoal the robot can move to. The implementation of grid-world can be found in https://github.com/mit-acl/dynus-vlm.

### v2

@@ Abstract
We present a novel high-level planning framework that leverages vision-language models (VLMs) to improve autonomous navigation in unknown indoor environments with many dead ends. Traditional exploration methods often take inefficient routes due to limited global reasoning and reliance on local heuristics. In contrast, our approach enables a VLM to reason directly about occupancy maps in a zero-shot manner, selecting subgoals that are likely to yield more efficient paths. At each planning step, we convert a 3D occupancy grid into a partial 2D map of the environment, and generate candidate subgoals. Each subgoal is then evaluated and ranked against other candidates by the model. We integrate this planning scheme into DYNUS [1], a state-of-the-art trajectory planner, and demonstrate improved navigation efficiency in simulation. The VLM infers structural patterns (e.g., rooms, corridors) from incomplete maps and balances the need to make progress toward a goal against the risk of entering unknown space. This reduces common greedy failures (e.g., detouring into small rooms) and achieves about 10% shorter paths on average.
@@ I Introduction
Autonomous mobile robots can be applied in numerous domains such as package delivery and search-and-rescue. In many of these applications, the robots need to operate in dense, cluttered, unknown environments where prior maps are not available. When navigating such spaces, traditional approaches such as DYNUS [1] often take inefficient routes by entering dead ends like small rooms, due to their lack of higher-level reasoning abilities. For instance, DYNUS simply projects the goal location onto a local map around the robot, and generates paths without global reasoning.
In recent years, many works have been proposed to improve the planning capabilities of robotic systems by employing large language models (LLMs) [2, 3, 4], including in frontier exploration schemes [5] and object-goal navigation [6]. We aim to leverage the reasoning competencies of multi-modal LLMs, in particular VLMs, to achieve efficient routing in unknown indoor environments that typically contain many impasses. The novelty of our approach lies in allowing VLMs to directly reason about structural cues from incomplete maps, rather than relying on RGB images. At each step, we convert the 3D occupancy grid into a voxelized 2D map, generate candidate subgoals, and ask the VLM to score directions in the context of the partial map (e.g., infer likely rooms/corridors from partial structure despite occlusions). We then aggregate multiple VLM queries, rank candidates by expected path efficiency, and select the top subgoal.
@@ III-A Problem Formulation
Our goal is to enable autonomous robots to efficiently reach a specified target in initially unknown indoor environments, while avoiding dead ends that require backtracking. We formulate this as a goal-directed exploration problem under partial observability. At each decision step, the high-level planner receives (1) the terminal goal position and (2) a 3D occupancy grid that is built online, and outputs a navigation subgoal that is likely to result in the most direct possible path to the goal.
@@ III-B Representing the Environment
The only information provided by the robot is a 3D occupancy grid of the environment. We do not rely on RGB images, which are constrained by lighting conditions and limited fields of view, but instead use occupancy data that capture the environment’s structure holistically and enable the system to operate effectively even in low-light conditions. However, the vision-language model cannot parse and understand raw sensor data or a voxel map. To obtain useful navigation guidance from the VLM, we ground the VLM with a map-like representation: a projected 2D occupancy map annotated with the robot pose, the goal, and candidate subgoals. This makes spatial relations explicit so the VLM can reason about where to go next. An example of such a representation is shown in Fig. 2.
@@ III-C Converting a 3D Occupancy Grid to a 2D Map
We keep a voxel map updated online. The map encodes unknown/free/occupied cells and overlays the robot pose, the goal (or its projection), and candidate subgoals. A cell is marked as free only if there is sufficient $z$-axis clearance for the robot and is reachable. We pass the map to the model either as a string or as an image. This can be achieved by associating each grid cell with a character or a color, respectively. Because the input grid resolution is fixed, we specify the desired size of the map in both meters and grid cells, and down-sample the grid accordingly. To get the best possible approximation:
@@ III-F Selecting a Subgoal
The median has the advantage that it is robust to outliers, and does not blend different probabilities. For instance, taking the mean of two valid but different solutions may not necessarily yield a good answer. On top of that, we compute the median absolute deviation (MAD) for each belief as an uncertainty measure:
@@ V Conclusions
This paper presents a VLM-based high-level planning framework that selects navigation subgoals at frontiers between known and unknown space in indoor environments. By providing occupancy maps as input, the model can reason about basic geometric structure (e.g., rooms and corridors). The method assumes minimal prior knowledge about the environment or sensors, and integrates cleanly with existing autonomy stacks.
In partially observed settings, reaching a goal while avoiding unnecessary detours or dead ends requires making decisions under uncertainty. Our simulations show that the proposed approach avoids most dead ends and yields shorter paths on average than the baseline in DYNUS [1]. We also observe that VLMs can infer structural patterns from incomplete maps and make reasonable trade-offs when uncertainty is high.
The main limitation is model latency, which constrains real-time deployment. We expect this to improve with faster hardware and models. Future work could explore incorporating region-level semantic information to further inform decisions and support richer trade-offs.
Supplementary Material
V-A Code
The implementation of our algorithm can be found at https://github.com/mit-acl/dynus-vlm.
V-B grid-world Simulation Tool
See Fig. S1.
Figure S1: View of grid-world, a simulation tool we built to study the navigation capabilities of LLMs/VLMs. The 2D map on the left shows the robot (red) traversing a corridor in an effort to reach the final goal (green) in the top right. The squares in the middle and on the right display the normalized median belief and the median absolute deviation (MAD) of the belief, respectively, for each candidate subgoal the robot can move to. The implementation of grid-world is also included in the provided code.
V-C Prompt
@@ 1. Candidate Move Evaluation
Developer Prompt (2/2) - Enclosure Check:
@@
- Progress Toward Goal:
@@
2. Candidate Move Comparison
@@ V-D Example Responses
    “This candidate is to the far north, up an extended corridor and within a semi-enclosed room. There is no visible path southward toward the projected goal. The route is circuitous and likely inefficient, with higher enclosure risk.”
    “Southernmost exit from central area, leading south and slightly east. Appears to open onto a possible corridor or hallway extending south, roughly matching the direction of the goal. No signs of imminent enclosure.”
@@
    “Northeast frontier, but accessing it requires backtracking through already-explored space. No clear indication it leads toward the goal. Likely a dead-end room given the local geometry and surrounding walls.”
    “This candidate is located at the top-left part of the explored area, apparently in a cul-de-sac (enclosed by walls on north and west sides). No visible path toward the goal projection to the far southeast. High risk of being a dead end.”
@@
    “This candidate is north of the robot and to the east, leading into a corridor in a section with more open unexplored area. There’s a good chance it opens up, but it’s not the closest to the direction of the goal projection.”
    “This is the southernmost candidate, ending the current visible main corridor. It is the frontier closest to the projected goal position, opens directly onto unknown space, and is highly promising for discovering the route to the goal. However, possible enclosure just south, so there is still risk.”
@@
    “This is to the south of the current room and could be a passageway, but still not fully aligned with the more southeast position of the goal. Low enclosure risk, but the area seems more like an open room than a corridor.”
    “This move goes south, away from the central room, and may be a corridor leading farther south. Though not directly leading southeast, it’s the only exit on the lower half; if this is a corridor, it could bend toward the east, following the general direction of the goal projection. Medium uncertainty.”